An alpha-particle has a charge of +2e and a mass o
Answer #1

Data Given Charge on the alpha particle q = 3.2\times 10^{-19} kg , Mass of the alpha particle m= 6.64\times 10^{-27} C

Potential Difference V= 1.10\times 10^{6} V And Magnetic field B = 2.80 T

(a) For the speed of alpha particle

We know that P.D = \frac{Wrok Done}{Charge}\Rightarrow Work Done = P.D\times Charge

From work energy theorem Work done = Change in K.E

So K.E. = \frac{1}{2}m.v^{2} = P.D\times q

v^{2} = \frac{P.D\times q}{m}= \frac{2\times 1.1\times 10^{6}\times 3.2\times 10^{-19}}{6.64\times 10^{-27}}v^{2} = \frac{P.D\times q}{m}= \frac{2\times 1.1\times 10^{6}\times 3.2\times 10^{-19}}{6.64\times 10^{-27}}= 1.06\times 10^{14}

v = \sqrt{1.06\times 10^{14}} = 1.03\times 10^{7} m/s

(b)

Magnetic Force on the alpha particle is given by

F_{B} = Bqv\sin \theta = B.q.v= 2.80\times 3.2\times 10^{-19}\times 1.03\times 10^{7}= 9.22\times 10^{-12} N

(c) Radius of Circular Path is given by

R = \frac{m.v}{B.q}= \frac{6.64\times 10^{-27}\times 1.03\times 10^{7}}{2.8\times 3.2\times 10^{-19}} =0.7633 \times 10^{1}= 7.633 m